OpenStudy (anonymous):

Solve the equation with the quadratic formula. r^2 + 9r - 22 = 0

OpenStudy (anonymous):

1) Put the quadratic equation into standard form. 2) Write out the formula and what a, b, & c stand for. 3) Substitute for each variable. 4) Split into two separate equations (setting each equal to zero) and solve. 5) Check each root in the original equation.

OpenStudy (anonymous):

r^2 +9r -22 = 0 (r + 22)(r -?)=0 (r+22)=(r-?) So something along these lines.

OpenStudy (anonymous):

r=2,−11